OpenStudy (anonymous):

sq. rt. of (3x - 2) = x - 2

OpenStudy (anonymous):

lol really?

OpenStudy (anonymous):

x^2 -4x +4 = 3x-2

OpenStudy (anonymous):

x^2 -7x +6

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

then... \[{x_{1/2}} = \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\]

OpenStudy (anonymous):

whats funny ?

OpenStudy (anonymous):

Good question @tana

OpenStudy (anonymous):

(x-6)(x-1)=0 ==> x=6 or x=1

OpenStudy (anonymous):

now answer my question... Are the Euler-Mascheroni constant...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

lol^

OpenStudy (anonymous):

nice one